This is the problem:
Write an equation of the line with the given slope that contains the given point.
Q(5,2),slope -2

Answer by tutorcecilia(2144)

Q(5,2),slope -2
Use the point-slope equation of a line because you are given a point and a slope]
.
(y-y1)=m(x-x1)
(y-2)=-2(x-5) [Plug-in the values]
y-2=-2x+10 [Isolate the y]
y=-2x+10+2
y=-2x+12